OpenStudy (anonymous):

HELP!!! Consider the equations I. v = v0 + ax II. y = (2m) cos(kx), where k = 2m^-1 III. v^2 = ax Which ones are dimensionally correct?

OpenStudy (anonymous):

I v = v0 + ax if x=time, correct II y = (2 m) cos(kx), where k = 2 m−1 y=2cos2x x=time, incorrect. III v2 = ax if x=time, incorrect

OpenStudy (anonymous):

That is what I thought at first, but I don't think it's right.

OpenStudy (anonymous):

why ? what the problem ?!

OpenStudy (anonymous):

In case x isn't = time, suppose x represents a distance.

OpenStudy (anonymous):

Ooh ! ya thats right because if we suppose x represents a distance ,, it`s become : v = v0 + ax if v is velocity, x is distance and a is acceleration, this is not correct !

OpenStudy (anonymous):

So, which ones would be correct? So, I know v = v0 + ax is not correct, but the other two?

OpenStudy (anonymous):

I think the second one is correct!

OpenStudy (anonymous):

no i don`t think. because if we try like that : the dimension of v is Length/Time while the dimension of (a.x) is L²/T² ..we find that one is not correct ! So ,,maybe the third one is correct : v² = ax [v²] = (L / T)² = L² / T² [ax] =[a][x] = (L / T²) L = L² / T² therefore [v²] = [ax] (L and T stand for Length and Time dimensions)

OpenStudy (anonymous):

So, just the third one will be correct?

OpenStudy (anonymous):

yes .. just the third one

OpenStudy (anonymous):

Because the second one is y = (2m) cos(kx), where k = 2m^-1.

OpenStudy (anonymous):

Equation II is not correct because [y] = Length (y is a displacement in harmonic oscillation) while both 2π and cos(2π-1)x are dimensionless.

OpenStudy (anonymous):

Oh yeah! That's right! So, just Equation III will be the correct one, right?

OpenStudy (anonymous):

right :)

OpenStudy (anonymous):

Okay, thank you so much! I appreciate it! :-)

OpenStudy (anonymous):

Welcome :)